OpenStudy (anonymous):

SOMEONE PLEASE HELP ME SO I CAN GO TO BED. match with their equivalent: y = -1/2x + 2 x - 2y = 4 y - 3 = 1/2(x - 4) y = -1/2x + 4 y = 1/2x + 9 ----------------------- y - 5 =1/2(x + 8) x + 2y = 4 x + 2y = 8 y =1/2x + 1 y =1/2x - 2

OpenStudy (anonymous):

Okay so for the bottom half what you want to do is distribute and you'll get the same answer that is on top

OpenStudy (anonymous):

So for y-5=1/2(x+8) Distribute Y-5=1/2x+4 Then add 5 to both sides to get y by itself Y=1/2x+9

OpenStudy (anonymous):

Then the next one try to get y by itself X+2y=4 Subtract x to both sides 2y=-x+4 Divide by 2 to get y by itself Y=-1/2x+2

OpenStudy (anonymous):

I hope you get the idea and hope this helps. If it does give me a medal thanks 😄

OpenStudy (anonymous):

marry me?? :DDDD